Circular No. 6947

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 M1 (LINEAR) Parabolic orbital elements from MPEC 1998-M05 (which included pre-

discovery May 22 LINEAR observations):

T = 1998 Oct. 29.2268 TT  ω = 19.^{o}8009
Ω = 256.0651 }2000.0
q = 3.110467 AU  i = 20.3890
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 16 15^{h}27.^{m}20 −25^{o}03.^{′}6 2.454 3.375 150.^{o}4 8.^{o}5 15.2
26 15 25.74 −23 46.9 2.491 3.339 140.6 11.1 15.2
July 6 15 26.06 −22 36.5 2.548 3.305 131.0 13.4 15.2
16 15 28.29 −21 34.8 2.623 3.274 121.8 15.3 15.2
26 15 32.43 −20 43.1 2.711 3.245 113.0 16.7 15.3

On June 17 and 18 UT, several CCD observers, including W. Offutt (Cloudcroft, NM) and C. Hergenrother (Mount Hopkins), reported a general asymmetrical coma and fan tail extending ∼ 20′′ in p.a. 150^{o}–160^{o}, then curving west-southwest toward p.a. 250^{o}, where it was ∼ 15′′–20′′ in width on June 17.15 (Offutt), continuing for over 6′ (Hergenrother).

SUPERNOVA 1998cn IN NGC 3735 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1998cn (cf. IAUC 6944), obtained by B. Carter on June 19.2 UT with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows the supernova to be of type Ia, ∼ 3 weeks past maximum light. The host galaxy’s recession velocity, as determined by narrow emission lines, is 2800 km/s.

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates: May 18.47 UT, 8.4 (D. A. J. Seargent, Cowra,

N.S.W., 10×50 binoculars); 23.35, 9.3 (A. Hale, Loomberah, N.S.W., 10×70 binoculars); 27.34, 8.5 (J. Seach, Chatsworth Island, N.S.W., 10×50 binoculars); June 3.36, 9.4 (Seach, 0.25-m reflector); 17.38, 9.8 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector).

1998 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green